A takes 5 days more than B to do a certain job and 9 days more than C; A and B together can do the job in
the same time as C. Then A alone can do the job in

Grade:12th Pass

1 Answers

Mayank Jain
13 years ago

Let the day taken by C alone be n

so in one day work done by C=1/n

Time taken by A to do the same work = n+9 days

so in one day work done by A=1/n+9

Also A and B can together do the work in same time as C

so work done in one day by A and B together = work done by C in one day=1/n

let work done by B in one day be b

so 1/n=b+(1/n+9)

so b=9/n(n+9)

 

Hence B takes n(n+9)/9 days to finish the work alone.

According to Question days taken by A alone = days taken by B alone +5

so n+9=n(n+9)/9+5

solving this n2=36 and n=6 days

Hence A alone can do the job in n+9=6+9=15 days time

 

Please approve if correct.